# Wigner’s friend scenarios: on what to condition and how to verify the predictions

### v1

@@ IV Measurement in Wigner’s bubble
Let us suppose that the random choice determines that measurement MW in which Wigner measures in the preferred basis of his environment $|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}\}$ should be made next. On the basis of his state assignment (5), Wigner makes a prediction that the two outcomes will occur with probabilities $p^{W}_{\omega_{0}}=1,p^{W}_{\omega_{1}}=0$. He then hands this prediction to referee RW. The aforementioned prediction is equally applicable also in the case when the SWAP operation is applied prior to the MW measurement.
@@
On the other hand, when applying the SWAP operation the state of the friend is just Eq. (6) where index F have been replaced by A. She predicts that probabilities for observing one or the other outcome of MW are $p^{F}_{\omega_{0}}=p^{F}_{\omega_{1}}=1/2$ (not to be confused with the friend’s probabilities of the first qubit measurement). These numbers can be encoded in a degree of freedom of her laboratory that did not get correlated to the outcome of the previous qubit measurement. Since the message containing the prediction also carries no information about the outcome of friend’s qubit measurement, it factorises out from her laboratory and can be transmitted to Wigner’s laboratory to the responsible referee RW as illustrated in Fig. 2.
@@
Figure 2: The schematic illustration of the game protocol in the case that measurement MW in Wigner’s bubble is chosen to be performed. On the basis of quantum state assignments in their respective laboratories, the two players make their predictions and submit them to the responsible referee RW in Wigner’s bubble. The probabilistic prediction of the friend contains no information about the outcome of the qubit observed in her laboratory. Consequently, its communication to the referee has no effect on the state of the laboratory as described by Wigner. The outcome of the MW measurement is in accordance with the prediction of Wigner, and in repeated trials, a discrepancy from the prediction of the friend can be observed.
@@
The measurement is then performed in Wigner’s bubble and the result $\omega_{0}$ is observed. Since the friend and Wigner predictions differ in the estimated probabilities, one would have to perform a sufficient number of runs to verify the difference between the two predictions with a certain degree of confidence, always keeping the records of both the measurement outcomes and the two predictions of the previous runs. The observed relative frequency will eventually show a significant deviation from the friend’s prediction as given e.g. by Wald’s sequential probability ratio test [31]  [^3] One can also construct a scenario in which Wigner’s assigns probability one to a specific outcome, while the friend assigns a probability that is arbitrarily close to zero, see Ref. [8]. This would render the argument arbitrarily close to a deterministic one. and consequently the referee RW will award Wigner. Interestingly, this situation is maintained even in the cases where partial information about the qubit measurement outcomes are leaked into Wigner’s environment, as explicitly discussed in Appendix B.
@@
Note that the measurement procedure affects the friend’s memory in that the content of the memory that contained the information about the outcome have been swapped with the ancilla state. While this allowed the friend to avoid the problem of self-reference, her memory of the outcome was lost in the process. One might ask, then, whether a partial loss of the memory would alter the entire game to such an extent that the friend’s identity as an observer would be lost, and she could no longer be regarded as the same player who started the game. We leave this point for the discussion, and remind the reader at this point that there is still a record (possibly with her signature) of the friend’s previous prediction, which she can inspect and accept as valid evidence of her previous prediction.
@@
The result of the game shows us that the friend could have made a more accurate prediction if she adopted the Wigner state assignment whenever the measurement MW is chosen. In fact, she can do this, since there are no fundamental restrictions as to why this information should be inaccessible to her. In fact, Wigner could simply send this information to her bubble, or the friend herself could first construct the experimental situation from Wigner’s perspective and then enter the sealed laboratory that becomes her bubble.
V Measurement in the friend’s bubble
@@
Next we consider the case where the decision is made to take the measurement MF. This means that the measurement of the qubit along the z-direction in the friend bubble using projectors $\Pi_{\uparrow}$ and $\Pi_{\downarrow}$ is repeated. The responsible referee is RF. The game is illustrated in Fig. 3.
@@
Figure 3: The schematic illustration of the game protocol in the case that measurement MF in the friend’s bubble is chosen to be performed. The two players make their predictions on the basis of their state assignments and submit them to the responsible referee RF in the friend’s bubble. The prediction of Wigner can be sent inside the friend’s laboratory with no additional effects. The outcome of the measurement is consistent with the prediction of the friend, and in repeated trials, a discrepancy between the observed relative frequencies and the prediction of Wigner can be observed.
@@
Which state is the friend going to use as a basis for her prediction? After the action of the unitary interaction $U_{\text{int}}$ the friend’s state becomes Eq. (6). It is important to note that this is not the state that the friend would normally adopt in order to make predictions about measurements in her laboratory. Rather, it is the state that the friend estimates she should use for predicting results of measurement in Wigner’s bubble, such as MW, after the action of the unitary and conditional on knowledge of her measurement outcome. In fact, as discussed below Eq. (7), from the perspective of the friend, the state of her memory (or of the ancilla after the SWAP) remains in a definite state perfectly correlated with that of the spin even after the action of the interaction unitary. For all subsequent measurements in her lab, all that matters for predicting future results is the last record in her lab. And that will be either definitely $\uparrow$ or definitely $\downarrow$ [^4] It may happen, however, that the action of the unitary induces a flip in the friend’s memory, but even then the friend will still have a definite record of the first measurement in her bubble. All her subsequent predictions can then be conditioned on that record. Alternatively, in “many-worlds” terms, one might say that unitary evolution “splits” the world into two branches, so that in each branch a friend sees a definite (and different) outcome. Also in this view, in order to predict the outcome of any future measurement within her branch, the friend should condition her prediction only on the last record in her branch.. Thus, the friend’s description of the relevant degrees of freedom for predicting future measurement results in her laboratory remains to be given by state (3), even after the action of the unitary interaction.
@@
Hence, if she repeats the same measurement, she will always confirm the record of the previous measurement result (which she had after the action of the unitary): she finds the outcome $\uparrow$ with probability $1$ and $\downarrow$ with probability 00 if the record of her previous observation was $\uparrow$, and similarly she finds $\downarrow$ with probability $1$ and $\uparrow$ with probability 00 if the record of her first measurement was $\downarrow$. Hence, she predicts $p^{F}_{\phi_{0}}=1$ (or 0), $p^{F}_{\phi_{1}}=0$ (or 1) and gives this prediction to referee RF. Formally, the state assigned by the friend after the repeated measurement is completed is given by:
@@
$|\Psi^{W}_{4}\rangle$ $=|\mathrm{GHZ}\rangle_{SFF^{\prime}}|\omega_{0}\rangle_{W}$
$=(\hat{U}_{\mathrm{meas}}\otimes\mathds{1}_{F})|\Phi^{+}\rangle_{SF}|\phi_{0}\rangle_{F^{\prime}}|\omega_{0}\rangle_{W},$ (11)
@@
We conclude that the friend’s state assignment was more accurate than Wigner’s for a measurement carried out in her laboratory. However, unlike the case of the MW measurement, where the friend could adopt Wigner’s prediction before the measurement was carried out and thus estimate the results of the measurement as well as Wigner, this is not true in the present case since any leakage of the information about the result of the first friend’s measurement to Wigner would correspondingly collapse his state.
VI (In)compatibility between different bubbles
@@
In the case where the information about which outcome the friend has observed cannot be unambiguously retrieved from the Wigner’s environment, Wigner and the friend give different, irreconcilable state assignments. None of them is “right” and the other “wrong”. They live in different “bubbles” [12] and each of the state assignment is “right” relative to the measurements performed in their respective bubbles. In these cases both Wigner and her friend might be tempted to think that the other party is making a wrong assignment. On the one hand, Wigner may think that the operation performed by the friend was not actually a “proper measurement”, since, following his description, unambiguous information about the outcomes is not detectable anywhere in his bubble. On the other hand, the friend –who has evidences that a proper measurement actually happened– may think that Wigner’s unitary description was incorrect, since it did not include the spread of the information about the outcomes actually generated in her measurement. Nevertheless, the assumption that no observer is preferred and that observers in other bubbles may experience definite outcomes, just as they do in their own, may prompt them to consider the possibility of bubbles with fundamentally incompatible outcomes.
@@
On the other hand, as explicitly shown in Appendix B, in the special case in which information about “which-outcome” is completely leaked into Wigner’s environment (i.e. each outcome is correlated to an accessible distinguishable state of the environment), the measurement update rule can be safely applied by both observers. One can say that if both observers share the information about the outcomes then they live in the same “bubble”. However, it is enough to have an arbitrarily small uncertainty in Wigner’s environmental states about the result of the qubit measurement in friend’s lab to create an (in principle) detectable incompatibility between Wigner and the friend’s state assignment. Interestingly, the number of runs needed to determine the validity of Wigner (friend) predictions versus the alternative one is a monotonically increasing function of the encoding parameter representing the amount of leaked information (see Appendix B): it increases as the amount of information leaked into Wigner’s environment increasing and diverges when the two descriptions are compatible.
@@
It is important to notice that the assumption usually adopted that the friend must be in a perfectly close lab is unwarranted, but the distinction lies in what kind of information is leaked outside: Any unsubstantial information about the measurement procedure, like e.g. the information that the system and the friend are entangled for Wigner, or about whether forecast, a film critique etc, Wigner and the friend can communicate without affecting their state assignments. In a similar manner, Wigner and his friend can communicate their predictions, share the random choice determining the subsequent measurement, or use communication to ascertain that their descriptions are incompatible.
VII Conclusions
The fundamental principles of textbook quantum mechanics indicate that an observer may apply two distinct forms of evolution to a given quantum state, contingent upon whether the evolution is between measurements or after the measurement. In the former case, a unitary transformation is applied, whereas in the latter, projections are employed (in the ideal case). However, the textbooks are silent on the conditions that specify which evolution should be applied in a concrete situation. This ambiguity can be attributed to a fundamental issue in defining the characteristics distinguishing between laboratory devices that serve as measuring instruments and those that implement unitaries – an issue known as “the quantum measurement problem”.
In this work, we do not address the measurement problem. Rather, we accept the fact that through everyday praxis, an observer acquires the experience of various devices and can sort them into those that serve as measurement apparata and those that implement unitary evolutions. This is the starting point for our discussion. Nevertheless, Wigner’s thought experiment and its modern extensions present an additional challenge to the consistency of the textbook quantum mechanics. This is because different observers may assign different quantum states to the same system. One possible way to overcome this challenge is to adopt the view that the fundamental concepts of quantum mechanics, such as quantum state, unitary evolution, or updated state, can only be defined relative to a bubble. Within a single bubble, all hypothetical observers may share their knowledge, thereby establishing intersubjectivity within the bubble. Since a bubble is defined only in terms of the (kind of) information that is in principle available in a certain region, we call this “relative objectivity”. In fact, according to this view there is in general not a universal correct state, but once the physical conditions are specified (that is, the bubble is defined) this fixes a state. This is reminiscent of relativity theory, where for example simultaneity between space-like separated events is not absolutely defined, but this relation is fixed within any chosen foliation. It would be desirable to find analogous ways to formally transform the description of a bubble into another by developing the analogue of Lorentz transformations.
Here, we examine scenarios in which observers are able to transition between bubbles. The observers make predictions about measurements that will be made on their own bubbles, including their own memory. After the measurement, they can end up in another bubble and check the accuracy of their previous predictions. Our findings necessitate a refinement of the concept of objectivity relative to the bubbles. We showed that in certain circumstances, observers are eligible to adopt the state assignment from other bubbles from which measurements are performed and make predictions accordingly. In fact, the friend can have information about her state as seen by Wigner without decohering her state, but informing Wigner about her spin state would collapse the state for him. While in this paper we consider the two players in a more symmetric role, this difference still differentiates between them.
@@ VII Conclusions
If we consider the possibility of a hierarchy of a large numbers of bubbles, each of which can be measured from the ones “higher” it in the hierarchy, we may conclude that a complete description of the physical situation of each observer would also require a hierarchy of joint state assignments: the friend’s state for measurements in the friend’s bubble, Wigner’s state for measurements in Wigner’s bubble, and so on. It is important to reiterate that an observer cannot have the state assignment of any bubbles “below” themselves in the hierarhy without decohering them. For instance, Wigner cannot have the state assignment of the friend without collapsing it. But it is this asymmetry that allows the “lower” observer in the hierarchy to be more successful in certain tasks than the “higher” level observer.

### v2

@@ IV.2 Measurement in Wigner’s bubble
Let us suppose that the random choice determines that the measurement $M_{W}$, in which Wigner measures in the preferred basis of his environment $\{|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}\}$, should be made next. On the basis of his state assignment (5), Wigner makes a prediction that the two outcomes will occur with probabilities $p^{W}_{\omega_{0}}=1,p^{W}_{\omega_{1}}=0$. He then hands this prediction to referee RW. The aforementioned prediction is equally applicable also in the case when the SWAP operation is applied prior to the $M_{W}$ measurement.
@@
On the other hand, when applying the SWAP operation the state of the friend is just Eq. (6) where index F have been replaced by A. She predicts that probabilities for observing one or the other outcome of $M_{W}$ are $p^{F}_{\omega_{0}}=p^{F}_{\omega_{1}}=1/2$ (not to be confused with the friend’s probabilities of the first qubit measurement). These predicted values for the probabilities can be encoded in a degree of freedom of her laboratory that did not get correlated to the outcome of the previous qubit measurement. Since the message containing the prediction also carries no information about the outcome of friend’s qubit measurement, it factorises out from her laboratory and can be transmitted to Wigner’s laboratory to the responsible referee RW as illustrated in Fig. 2.
@@
Figure 2: The schematic illustration of the game protocol in the case that measurement $M_{W}$ in Wigner’s bubble is chosen to be performed. On the basis of quantum state assignments in their respective laboratories, the two players make their predictions and submit them to the responsible referee RW in Wigner’s bubble. The probabilistic prediction of the friend contains no information about the outcome of the qubit observed in her laboratory. Consequently, its communication to the referee has no effect on the state of the laboratory as described by Wigner. The outcome of the $M_{W}$ measurement is in accordance with the prediction of Wigner, and in repeated trials, a discrepancy from the prediction of the friend can be observed.
@@
The measurement is then performed in Wigner’s bubble and the result $\omega_{0}$ is observed. Since the friend and Wigner predictions differ in the estimated probabilities, one would have to perform a sufficient number of runs to verify the difference between the two predictions with a certain degree of confidence, always keeping the records of both the measurement outcomes and the two predictions of the previous runs. The observed relative frequency will eventually show a significant deviation from the friend’s prediction as given, e.g., by Wald’s sequential probability ratio test [34]. [^4] One can also construct a scenario in which Wigner’s assigns probability one to a specific outcome, while the friend assigns a probability that is arbitrarily close to zero, see Ref. [8]. This would render the argument arbitrarily close to a deterministic one. and consequently the referee RW will award Wigner. Interestingly, this situation is maintained even in the cases where partial information about the qubit measurement outcomes are leaked into Wigner’s environment, as explicitly discussed in Appendix B.
@@
Note that the measurement procedure affects the friend’s memory in that the content of the memory that contained the information about the outcome have been swapped with the ancilla state. While this allowed the friend to avoid the problem of self-reference, her memory of the outcome was lost in the process. One might ask, then, whether the loss of this part of the memory would alter the entire game to such an extent that the friend’s identity as an observer would be lost, and she could no longer be regarded as the same player who started the game. We leave this point for the discussion, and remind the reader at this point that there is still a record (possibly with her signature) of the friend’s previous prediction, which she can inspect and accept as valid evidence of her previous prediction.
@@
The result of the game shows us that the friend could have made a more accurate prediction if she adopted the Wigner state assignment whenever the measurement $M_{W}$ is chosen. In fact, she can do this, since there are no fundamental restrictions as to why this information should be inaccessible to her. In fact, Wigner could simply send this information to her bubble, or the friend herself could first construct the experimental situation from Wigner’s perspective and then enter the sealed laboratory that becomes her bubble.
IV.3 Measurement in the friend’s bubble
@@
Next we consider the case where the decision is made to take the measurement $M_{F}$. This means that the measurement of the qubit along the $z$-direction in the friend bubble using projectors $\Pi_{\uparrow}$ and $\Pi_{\downarrow}$ is repeated. The responsible referee is RF. The game is illustrated in Fig. 3.
@@
Figure 3: The schematic illustration of the game protocol in the case that measurement $M_{F}$ in the friend’s bubble is chosen to be performed. The two players make their predictions on the basis of their state assignments and submit them to the responsible referee RF in the friend’s bubble. The prediction of Wigner can be sent inside the friend’s laboratory with no additional effects. The outcome of the measurement is consistent with the prediction of the friend, and in repeated trials, a discrepancy between the observed relative frequencies and the prediction of Wigner can be observed.
@@
Which state is the friend going to use as a basis for her prediction? After the action of the unitary interaction $U_{\text{int}}$ the friend’s state becomes Eq. (6). It is important to note that this is not the state that the friend would normally adopt in order to make predictions about measurements in her laboratory. Rather, it is the state that the friend estimates she should use for predicting results of measurement in Wigner’s bubble, such as $M_{W}$, after the action of the unitary and conditional on knowledge of her measurement outcome. In fact, as discussed below Eq. (7), from the perspective of the friend, the state of her memory (or of the ancilla after the SWAP) remains in a definite state perfectly correlated with that of the spin even after the action of the interaction unitary. For all subsequent measurements in her lab, all that matters for predicting future results is the last record in her lab. And that will be either definitely $\uparrow$ or definitely $\downarrow$. [^5] It may happen, however, that the action of the unitary induces a flip in the friend’s memory, but even then the friend will still have a definite record of the first measurement in her bubble. All her subsequent predictions can then be conditioned on that record. Alternatively, in “many-worlds” terms, one might say that unitary evolution “splits” the world into two branches (as given in each of the two lines of Eq. (7)), so that in each branch a friend sees a definite (and different) outcome. Also in this view, in order to predict the outcome of any future measurement within her branch, the friend should condition her prediction only on the last record in her branch. Thus, the friend’s description of the relevant degrees of freedom for predicting future measurement results in her laboratory remains to be given by state (3), even after the action of the unitary interaction.
@@
Hence, if she repeats the same measurement, she will always confirm the record of the previous measurement result (which she had after the action of the unitary): she finds the outcome $\uparrow$ with probability $1$ and $\downarrow$ with probability 00 if the record of her previous observation was $\uparrow$, and similarly she finds $\downarrow$ with probability $1$ and $\uparrow$ with probability 00 if the record of her first measurement was $\downarrow$. Hence, she predicts $p^{F}_{\phi_{0}}=1$ (or 0), $p^{F}_{\phi_{1}}=0$ (or 1) and gives this prediction to referee RF. Formally, the state assigned by the friend after the repeated measurement is given by:
@@
$|\Psi^{W}_{4}\rangle$ $=(\hat{U}_{\mathrm{meas}}\otimes\mathds{1}_{F})|\Phi^{+}\rangle_{SF}|\phi_{0}\rangle_{F^{\prime}}|\omega_{0}\rangle_{W}$ (11)
$=|\mathrm{GHZ}\rangle_{SFF^{\prime}}|\omega_{0}\rangle_{W},$
@@
We conclude that the friend’s state assignment was more accurate than Wigner’s for a measurement carried out in her laboratory. However, unlike the case of the $M_{W}$ measurement, where the friend could adopt Wigner’s prediction before the measurement was carried out and thus estimate the results of the measurement as well as Wigner, this is not true in the present case since any leakage of the information about the result of the first friend’s measurement to Wigner would correspondingly collapse his state.
V (In)compatibility between different bubbles
@@
In the case where the information about which outcome the friend has observed cannot be unambiguously retrieved from the Wigner’s environment, Wigner and the friend give different, irreconcilable state assignments. None of them is “right” and the other “wrong”. They live in different “bubbles” [12] and each of the state assignment is “right” relative to the measurements performed in their respective bubbles. In these cases both Wigner and her friend might be tempted to think that the other party is making a wrong assignment. On the one hand, Wigner may think that the operation performed by the friend was not actually a “proper measurement”, since, following his description, unambiguous information about the outcomes is not detectable anywhere in his bubble. On the other hand, the friend—having evidence that a genuine measurement has taken place—may regard Wigner’s unitary description as inadequate, since it fails to account for the for-all-practical-purposes irreversible imprint of the measurement outcome within her bubble. In any case, the assumption that no observer is preferred and that observers in other bubbles may experience definite outcomes, just as they do in their own, may prompt them to consider the possibility of bubbles with fundamentally incompatible outcomes.
@@
On the contrary, as explicitly shown in Appendix B, in the special case in which information about “which-outcome” is completely leaked into Wigner’s environment (i.e. each outcome is correlated to an accessible distinguishable state of the environment), the measurement update rule can be safely applied by both observers. One can say that if both observers share the information about the outcomes then they live in the same “bubble”. However, it is enough to have an arbitrarily small uncertainty in Wigner’s environmental states about the result of the qubit measurement in friend’s lab to create an (in principle) detectable incompatibility between Wigner and the friend’s state assignment. Interestingly, the number of runs needed to determine the validity of Wigner (friend) predictions versus the alternative one is a monotonically increasing function of the encoding parameter representing the amount of leaked information (see Appendix B): it increases as the amount of information leaked into Wigner’s environment increases and diverges when the two descriptions are compatible.
@@
It is important to note that the common assumption that the friend must be in a perfectly closed lab is unwarranted. The key distinction lies in the type of information that leaks outside: any non-substantial information about the measurement procedure—such as the fact that the system and the friend are entangled from Wigner’s perspective—or unrelated topics like the weather forecast, a film critique, etc., can be communicated between Wigner and the friend without affecting their state assignments. Similarly, Wigner and his friend can exchange their predictions, share the random choice determining the subsequent measurement, or use communication to verify that their descriptions are incompatible.
@@
Finally, one can consider the possibility of a hierarchy of numerous bubbles, each of which can be measured from the “higher” ones in the hierarchy. In that case, we may conclude that a complete description of each observer’s physical situation would require a corresponding hierarchy of joint state assignments: the friend’s state for measurements in the friend’s bubble, Wigner’s state for measurements in Wigner’s bubble, and so on. It is important to reiterate that an observer cannot have the state assignment of any bubbles “below” themselves in the hierarchy without decohering them. For instance, Wigner cannot have the state assignment of the friend without collapsing it. But it is this asymmetry that allows the “lower” observer in the hierarchy to be more successful in certain tasks than the “higher” level observer.
VI Conclusions
According to the fundamental principles of textbook quantum mechanics, an observer may apply two distinct types of evolution to a quantum state, contingent upon whether the evolution takes place between measurements or following a measurement. In the former case, a unitary transformation is applied, whereas in the latter, projections are employed (in the ideal case). However, textbooks are silent on the conditions that specify which evolution should be applied in a concrete situation. This ambiguity can be attributed to a fundamental issue in defining the characteristics distinguishing between laboratory devices that serve as measuring instruments and those that implement unitaries – an issue known as “the quantum measurement problem”.
In this work, we did not address the measurement problem. Rather, we acknowledge that through everyday praxis, an observer gains experience with various devices and can distinguish between those that function as measurement apparatuses and those that implement unitary evolutions. This is the starting point for our discussion. Nevertheless, Wigner’s thought experiment and its modern extensions present an additional challenge to the consistency of the textbook quantum mechanics. This is because different observers may assign different quantum states to the same system. One possible way to overcome this challenge is to adopt the view that the fundamental concepts of quantum mechanics, such as quantum state, unitary evolution, or updated state, can only be defined relative to a bubble. Within a single bubble, all hypothetical observers may share their knowledge, thereby establishing intersubjectivity within the bubble. Since a bubble is defined only in terms of the (kind of) information that is in principle available in a certain region, we call this “relative objectivity”. In fact, according to this view, there is generally no universally correct state. However, once the physical conditions are specified (i.e., the bubble is defined), the state is determined. This is reminiscent of relativity theory, where for example simultaneity between space-like separated events is not absolutely defined, but this relation is fixed within any chosen foliation [35]. It would be desirable to find analogous ways to formally transform the description of a bubble into another by developing the analogue of Lorentz transformations.
However, we believe that a transformation from one observer’s perspective to another—such as from the friend to Wigner—may not, in general, be representable by a unitary operation (as in the framework of quantum reference frame transformations [36]). In fact, from the friend’s perspective, the measurement has taken place, resulting in two or more classically distinguishable outcomes—i.e., distinct states corresponding to the possible measurement results. From Wigner’s perspective, however, the friend and the measured system remain in a single, entangled state. Crucially, since unitary transformations are bijective, they cannot map multiple classical states corresponding to the friend’s possible distinct experiences onto a single entangled state as described by Wigner. This suggests that no single unitary can bridge the two perspectives.
We examined scenarios in which observers can transition between bubbles, gaining (partial) knowledge of the state of another bubble. The observers are then tasked with making predictions about measurements that will be made in or on their own bubbles, including their own memory. Our findings highlight the need for a refinement of the concept of objectivity relative to the bubbles. We demonstrated that, under certain circumstances, observers are entitled to adopt the state assignment from other bubbles where measurements are performed and make predictions accordingly. A true novelty is that, in different cases, both Wigner and the friend can achieve this. This symmetrizes the scenario and challenges the view that Wigner, as a superobserver, always possesses more predictive power than the friend.